OpenStudy (anonymous):

just need my answer checked... will give an award! Question: 8(2n-5) =4n+ 8 My Answer: 6 *If got wrong please explain to me why! :) I now know I got it wrong... just help me figure it out please!

OpenStudy (anonymous):

With 6 entered in as n the problem would go as follows \[8\left( 2 * 6 - 5 \right) = \left( 4*6 \right) + 8\] \[88 = 32 \] which is false

OpenStudy (anonymous):

I so lost! I dont know how else to do it

OpenStudy (ryan123345):

Empty set

OpenStudy (anonymous):

I think the answer might be 4 I'll try entering that in

OpenStudy (anonymous):

\[8\left( 2*4-5 \right) = 4* 4 + 8\] \[8(3) = 16+ 8\] \[24 = 24\]

OpenStudy (anonymous):

8(2n-5) =4n+ 8 16n-40=4n+8 Add like terms. In this case you would subtract 4n from both sides and add 40 to both sides. 16n-4n=8+40 12n=48 Now divide both sides by 12. n=4

OpenStudy (anonymous):

yeah you are right its 4 Thanks.. but I dont get it if you are trying to straight up solve it..

OpenStudy (anonymous):

both.. never mind.. I got it

OpenStudy (anonymous):

@Jazmyne25 She explains it :)

OpenStudy (anonymous):

you and the other girl got a medal! :) Thanks so much for your help!

OpenStudy (anonymous):

No problem, anytime :)